OpenStudy (anonymous):

Find the value of 7 + 3^2 (-5 + 1) ÷ 2.

OpenStudy (anonymous):

4 i think. I'm about 173% sure tho

jigglypuff314 (jigglypuff314):

PEMDAS -5+1 = -4 3^2 = 9 9 times -4 = -36 -36 / 2 = -18 7 + (-18) = -11 Nurali was right :P